A room contains 48 kg of air. How many KWh of energy are necessary to heat the air in the house from 7oC to 28oC? The heat capacity of air is 1.03 J/goC.
1
Expert's answer
2015-10-19T05:09:41-0400
Q = C*m*(T2 - T1) = 1.03 J/(g*K)*48 kg*(28 C - 7 C) = 1.03 J/(g*K)*48000 g*(28 C - 7 C) = 1038240 J = 288.4 Wh = 0.2884 kWh

1 Wh = 3600 J